OpenStudy (anonymous):

Help me with absolute value equations, inequalities? I need to know every thing. HELP!

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

I don't care, really. I just need to know the reasoning, what it solves, why you do it etc. Here's an example: l 2x-4 l = 10

OpenStudy (anonymous):

first: absolute values can take on two different values, positive or negative, and returns a positive.

OpenStudy (anonymous):

for example: \[|-1| = 1\] and \[|1| = 1\] that's the basics of absolute values. now to more interesting stuff like your example.

OpenStudy (anonymous):

because it has 2 possible values, you have to break it down into 2 parts like this \[2x - 4 = 10\] and \[2x - 4 = -10\] from there, find the value of x.